Baka Corporation applies manufacturing overhead on the basis of direct labor-hours. At the beginning of the most recent year, the company based its predetermined overhead rate on total estimated overhead of $244,500 and 9,500 estimated direct labor-hours. Actual manufacturing overhead for the year amounted to $245,200 and actual direct labor-hours were 6,200.

The overhead for the year was: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $85,612 overapplied

  • $84,912 underapplied

  • $85,612 underapplied

  • $84,912 overapplied
Answer #1

Answer) $85612 underapplied

Explanation:

Predetermined Overhead rate = 244500 / 9500 = 25.74

Manufacturing Overhead applied to production = 6200 x 25.74 = 159588

Underapplied Overhead = Actual Overhead - Overhead applied to Production

= 245200 - 159588

= $ 85612